OpenStudy (yayrocks23):

PLEASE HELP ASAP When solving this equation by completing the square, which of the following shows the equation with the square completed? x2 - 4x - 32 = 0 I'm having a lot of trouble with quadratics :/

OpenStudy (bmk614):

ax2-bx+c c is b/2 and then squared

OpenStudy (bmk614):

so C would be 4

OpenStudy (yayrocks23):

I'm confused. The answer choices are x2 - 4x - 32 = 0 A)(x - 4)2 = 36 B)(x - 2)2 = 36 C)(x - 8)2 = 4 @bmk614

OpenStudy (bmk614):

well when you take that you need to add 32 to both sides but then add 4 to both sides. Then you factor the (x2-4x+4) it would be (x-2)2=36

OpenStudy (mathmale):

"equation with the square completed?" Please express your exponentiation with the symbol ^: x^2 - 4x - 32 = 0 correct x2 - 4x - 32 = 0 incorrect Given x^2 - 4x - 32 = 0, our task is to rewrite x^2-4x as a perfect square of a binomial. 1. Take the coefficient of the x (first power of x): Here it's -4. 2. Take half of that coefficient. Here, that would be -4/2, or -2. 3. Square this last result. Here, that would be (-2)^2, or 4. 4. Write x^2 - 4x - 32 = 0 as x2 - 4x - 32 = 0 Add that 4, and then subtract 4 (net is zero) in the empty space above. See if you can finish this work. Show all work you do, please. What is the square root of x^2-4x + 4?

OpenStudy (yayrocks23):

Thank you @bmk614 and @mathmale